OpenStudy (anonymous):

Can someone show me how to solve: lim xcos(1/x) x--> 0

OpenStudy (anonymous):

cosine is stuck between \(-1\) and \(1\) but \(x\to 0\)

OpenStudy (anonymous):

so the whole thing must go to zero cosine can't mess it up any, as it is bounded above and below

OpenStudy (loser66):

@bobobobobb can you get him? me not

OpenStudy (anonymous):

Yes @satellite73 can you explain it ?

OpenStudy (anonymous):

can i explain what?

OpenStudy (anonymous):

i understand it now

OpenStudy (anonymous):

i suppose your math teacher would be happier if you used the "squeeze" theorem \[-1\leq \cos(\frac{1}{x})\leq 1\] so \[-x\leq x\cos(\frac{1}{x})\leq x\]

OpenStudy (anonymous):

the limit on the left and right is clearly 0 so the limit in the middle must be 0 as well

OpenStudy (anonymous):

thanks:)